Can someone help me solve this?
Solve the system by the elimination method.
3x-5y=1
6x-10y=4

Answer by stanbon(75887)

Solve the system by the elimination method.
3x-5y=1
6x-10y=4
-----
Multiply thru the top equation by 2 to get:
6x -10y = 2
6x -10y = 4
------
Assuming the x's are thesame and te y's are the same you get:
0 - 0 = 2
0 = 2
------
Conclusion: The variables are not the same in the 2 equations.
There is no (x,y) that satisfies the system.
Cheers,
Stan H.
================